My teacher would like us to answer her question & her question is what is wrong with this quadratic equation which is 52-x2+x=10 and the 2 on beside the x should be an exponent!!

Answer by Edwin McCravy(20054)

What is wrong with:

   52 - x² + x = 10  ?

First of all it doesn't have 0 on the right side.
So we add -10 to both sides:

   42 - x² + x = 0

Secondly it is not in descending order,
so we write the x² term first, then the
x-term and finally the constant

  -x² + x + 42 = 0

Thirdly the leading coefficient (the coefficient of the
x² term) should be positive, so we multiply it
through by -1, changing all the signs

   x² - x - 42 = 0

Now it's easy to factor as

(x - 7)(x + 6) = 0

and solve with the zero-factor property:

x - 7 = 0;  x + 6 = 0
    x = 7;      x = -6

Edwin